OpenStudy (anonymous):

two similar polygons have perimeters of 28 and 40; what is the ratio of their areas? what is the ratio of their volumes? thank you :)

OpenStudy (raden):

28/40 = 7/10 therefore, the ratio of their areas = (7/10)^2 = 49 : 100 and the ratio of their volumes = (7/10)^3 = 343 : 1000

OpenStudy (anonymous):

wouldnt they all just simplify to 7/10 though?

OpenStudy (anonymous):

never mind, haha thank you!

OpenStudy (raden):

you're welcome ;)